# Multi–charged geometries with cosmological constant

### v1

@@ C.2 Addition of charge
Looking at the metric (C.1), we observe that its radial and sphere components contain information about $(a_{1},a_{2},p$, while their ratio does not:
@@
Note that this condition doesn’t fully fix the gauge: while $C$ is a scalar under repameterizations of $x$, function $R$ is not. Constructing the scalar quantity,
@@
and functions $(f_{1},f_{2},f_{3},f_{4})$ can be expanded in the power series in $q^{2}$. In the first order in $q^{2}$, one finds a solution that depends on seven integration constants, but all of them can be absorbed by diffeomorphisms and shifts the scalar field:
@@ C.3 Relations to known solutions
Let us now recover the solution which describes a four–dimensional black hole with a magnetic and dilatonic charge [53]. This geometry in the the coordinates defined in [53] is given by (3.4).

### v2

@@ C.2 Addition of charge
Looking at the metric (C.1), we observe that its radial and sphere components contain information about $(a_{1},a_{2},p)$, while their ratio does not:
@@
Note that this condition does not fully fix the gauge: while $C$ is a scalar under repameterizations of $x$, function $R$ is not. Constructing the scalar quantity,
@@
and functions $(f_{1},f_{2},f_{3},f_{4})$ can be expanded in the power series in $q^{2}$. In the first order in $q^{2}$, one finds a solution that depends on seven integration constants, but all of them can be absorbed by diffeomorphisms and shifts of the scalar field:
@@ C.3 Relations to known solutions
Let us now recover the solution which describes the four–dimensional black hole with a magnetic and dilatonic charge found in [121]. This geometry, in the coordinates defined in [121], is given by (3.4).